# Multi-Task Models Adversarial Attacks

### v3

@@ Abstract
Multi-Task Learning (MTL) involves developing a singular model, known as a multi-task model, to concurrently perform multiple tasks. While the security of single-task models has been thoroughly studied, multi-task models pose several critical security questions, such as 1) their vulnerability to single-task adversarial attacks, 2) the possibility of designing attacks that target multiple tasks, and 3) the impact of task sharing and adversarial training on their resilience to such attacks. This paper addresses these queries through detailed analysis and rigorous experimentation. First, we explore the adaptation of single-task white-box attacks to multi-task models and identify their limitations. We then introduce a novel attack framework, the Gradient Balancing Multi-Task Attack (GB-MTA), which treats attacking a multi-task model as an optimization problem. This problem, based on averaged relative loss change across tasks, is approximated as an integer linear programming problem. Extensive evaluations on MTL benchmarks, NYUv2 and Tiny-Taxonomy, demonstrate GB-MTA’s effectiveness against both standard and adversarially trained multi-task models. The results also highlight a trade-off between task accuracy improvement via parameter sharing and increased model vulnerability due to enhanced attack transferability.
Introduction
Multi-task learning (MTL) leverages a single machine learning model to simultaneously address multiple tasks, such as semantic segmentation, depth estimation, and other vision-based prediction tasks. The single model, called multi-task model, shares parameters between tasks and is shown to have lower inference costs and higher generalization performance compared to single-task models without parameter sharing (Ruder 2017; Zhang, Liu, and Guan 2022a; Yao et al. 2020). Multi-task models are widely used in practical applications with high-security requirements, such as robotics and autonomous driving (Leang et al. 2020; Kokkinos 2017; Arcari et al. 2023). In this paper, we focus on the branched multi-task models, which are the most representative in the MTL literature (Zhang, Liu, and Guan 2022b). We use the terms task sharing and parameter sharing interchangeably.
In parallel to the developments in MTL, the security of single task classifiers has been put into question due to the existence of adversarial examples (Goodfellow, Shlens, and Szegedy 2014). An adversarial example is an input to a machine learning model (typically an image) that has been manipulated such that the model misclassifies the example with high confidence, but a human can still correctly recognize the input. Adversarial examples can be generated through white-box or black-box attacks depending on the assumed capabilities of the attacker (Tramer et al. 2020; Mahmood et al. 2021a). White-box attacks are generally considered more powerful (Carlini et al. 2019) because the attacker has access to the parameters and structure of the trained model.
@@ Introduction
Although adversarial attacks have been extensively studied on single-task models (Liu et al. 2016; Mahmood et al. 2022; Xu et al. 2022), related work on multi-task models is scarce. A pioneering study (Mao et al. 2020) pointed out that the adversarial robustness of deep neural networks increases as the number of tasks increases. MTA (Guo et al. 2020) tries to develop attacks in the MTL setting; however, the adversarial samples generated are task-specific and thus fail to attack all tasks simultaneously. Some other works (Gurulingan, Arani, and Zonooz 2021; Sobh et al. 2021) attempted to attack multi-task models by generating adversarial examples for each image while attacking one task at a time. Several critical security research questions (RQ) on MTL remain unclear:
@@
RQ1: How secure are multi-task models to conventional single task adversarial attacks?
@@
RQ2: Can adversarial attacks be designed to attack multiple tasks simultaneously?
RQ3: Does task sharing and adversarial training increase multi-task model robustness to adversarial attacks?
This paper answers the three questions through careful analysis and rigorous experimentation. To answer RQ1, we develop two naïve adaptations of single task white-box attacks for multi-task models, and analyze their inherent drawbacks. To answer RQ2, we propose a novel attack framework, GB-MTA (Gradient Balancing Multi-Task Attacker) to generate adversarial samples effective in attacking all tasks in a multi-task model. GB-MTA frames the problem of finding a unified attack perturbation in MTL as an optimization problem based on the averaged relative loss change (Sun et al. 2019; Zhang, Liu, and Guan 2022a) across tasks and solves the problem by approximating it as an Integer Linear Programming (ILP). To answer RQ3, we experiment with different levels of task sharing and demonstrate that there is a fundamental trade-off: Improving task accuracy and model efficiency through parameter sharing can increase the model’s vulnerability to adversarial attacks designated for these related tasks. We further explore the defense side of MTL, by adversarially training models with examples generated by GB-MTA. Our contributions to advancing the security of the field of MTL are summarized as follows.
Dynamic Gradient Balancing Multi-task Attack Framework - We formulate the MTL adversarial attack as an optimization problem with a specially designed multi-task objective function. To solve this optimization problem, we introduce a novel approach, GB-MTA that balances gradients from multiple tasks in a multi-task model when creating adversarial samples that work across all tasks.
Multi-Task Models Robustness Trade-off - We empirically demonstrate that task sharing can undermine model robustness due to increased attack transferability. In NYUv2, as the level of task sharing increases, the task attack transferability increases by , , and , respectively, when attacking segmentation, depth estimation, or normal prediction.
@@
Multi-Task Learning Adversarial Training - We incorporate adversarial examples into the training of multi-task models to defend against adversarial attacks. The robustness of the models improves markedly, measured by the decreased accuracy drop after attacking from to . When attacking these adversarially trained models, GB-MTA still outperforms baselines by up to .
Attack Framework
This section first discusses existing white-box attacks and our adversarial threat model. This section then shows how single task white-box attacks can be adapted to multi-task models. We formulate the GB-MTA framework in the next section.
Single Task White-Box Attacks
@@
In general, adversarial attacks can be formulated as follows (Madry et al. 2018). Let represent a clean input and its corresponding label. An attacker adds an adversarial perturbation to the input , to maximize the value of a loss function :
@@
where denotes the parameters of the trained model under attack, and represents the maximum amount the adversary can perturb the input according to a given -norm. For notational simplicity, we omit in our future derivations.
@@
Threat Model: In this paper, we focus on the untargeted white-box adversarial threat model (Carlini et al. 2019) as this represents one of the strongest and most widely used adversarial machine learning formulations (Carlini and Wagner 2017; Dong et al. 2018; Croce and Hein 2020b). In this setup, the attacker has knowledge of the model structure, trained model parameters and the corresponding loss function . In terms of bounds on the adversarial perturbation, we use one of the most widely used norms, , in line with previous works (Guo et al. 2020; Xu et al. 2022; Rathbun et al. 2022).
@@
Single Task Attacks: In the white-box setting, one of the most prevalent strategies for generating the adversarial perturbation is to maximize the loss function by following the gradient ascent direction. This was originally done with the Fast Gradient Sign Method (FGSM) attack proposed in (Goodfellow, Shlens, and Szegedy 2014). Since the advent of FGSM, numerous improvements to the attack have been proposed. Although enumerating all the improvements in FGSM is beyond the scope of this paper, several important attack updates are worth noting. Updated attacks include the Projected Gradient Descent (PGD) attack (Madry et al. 2018), which adds a randomized start and makes FGSM iterative. The Momentum Iterative Momentum (MIM) (Dong et al. 2018) adds momentum to the gradient ascent optimization. More recently, in APGD (Croce and Hein 2020b), an adaptive step size has been shown to be one of the most effective white-box attacks for a single task, even against adversarial trained models (Mahmood, Mahmood, and Van Dijk 2021).
Naïve Multi-Task Attacks
RQ1: How secure are multi-task models to conventional single task adversarial attacks? This subsection answers the question by presenting two strategies for adapting single task white-box attacks to the multi-task formulation of the problem. It then discusses their inherent flaws, which are further demonstrated empirically. We denote these two attacks as naïve multi-task attacks.
@@
In the case of a single task, untargeted white-box attack, an adversarial example can be generated (Madry et al. 2018) iteratively:
@@
where represents the perturbation function associated with a specific white-box attack, represents the loss function for a single task, is the magnitude of the perturbation added in the current iteration of the attack and . Lastly, is the projection operation (Croce and Hein 2020b) to bound the adversarial sample within a specified range. In MTL, each input is associated with a set of true labels for tasks . Each task has its own task-specific loss function .
@@
Single Attack - The first way in which multi-task models can be attacked is by focusing on only a single task’s gradient and ignoring the gradients of the rest tasks (Sobh et al. 2021; Gurulingan, Arani, and Zonooz 2021). For example, from Eq. 2, APGD (Croce and Hein 2020b) can be adapted to attack a single task :
@@
where is a hyperparameter in APGD that controls the influence of previous update steps on the current update step. is the objective function of the task being attacked. In the experiment, we use Single-x to represent performing Single attack on a specific task x.
@@
Total Attack - The second way in which single task attacks can be converted to attack multi-task models is through totaling all associated task loss functions , via summation. For example, Eq. 2 with single task Projected Gradient Descent (PGD) (Madry et al. 2018) can be adapted for Total-PGD:
@@
Naïve Attack Limitations: Both the Single and Total attacks come with significant drawbacks. The effectiveness of the Single attack is based on the following assumption: using one task’s attack direction will guarantee attack success on all other tasks. However, this assumption does not always hold, as the effectiveness of a Single attack is restricted due to the limited transferability of attack methods (Mahmood, Mahmood, and Van Dijk 2021). For example, in Figure 1, we show that when attacking the segmentation or the depth estimation task solely (Single-Segm or Single-Depth), the effectiveness of the attack on the normal prediction task (Norm ARP) is limited. The definition of ARP is elaborated in the Experiments Section.
Likewise, the effectiveness of the Total attack is based on an underlying assumption: for an adversarial example to work across all tasks, no one task’s gradient should dominate to avoid the limited attack transferability issue that Single faces. We denote it as the non-dominant magnitude assumption. However, the issue of gradient dominance is well recognized within the MTL literature and has garnered significant attention in the field of MTL optimizers (Yu et al. 2020; Navon et al. 2022). Empirically, we observe from Figure 1 that the Total attack exhibits a pattern similar to Single-Segm, indicating that the segmentation task dominates the gradient directions.
We also consider a modified version of Total where we take the sign of the gradients before the summation. In this way, the non-dominant magnitude assumption can be circumvented. We denote this attack as SignTotal. However, we empirically show that this attack is also not effective in Section Results on Attack Performance, as completely ignoring task’s gradient magnitudes also leads to a suboptimal attack.
@@
GB-MTA Framework
RQ2: Can adversarial attacks be designed to attack multiple tasks simultaneously? Dynamic Gradient Balancing Multi-task Attack (GB-MTA) builds on the success of existing single task adversarial attacks, while addressing the challenge in attacking multi-task models. GB-MTA accomplishes this by actively balancing the gradients across tasks, to derive an adversarial perturbation that is effective on all tasks. In this section, we first formulate a new attack optimization problem tailored for multi-task models. Since the problem is intractable, GB-MTA reformulates it to an Integer Linear Programming (ILP) problem, and then generates adversarial samples by solving the ILP problem.
Multi-Task Attack Optimization: We first reformulate the original single-task adversarial optimization introduced in Eq. 1 by decomposing . Here, represents the magnitude of the perturbation. represents the signed gradient direction vector, with values .
The above formulation is used to attack a single task. Attacking multiple tasks simultaneously in a multi-task model is fundamentally a multi-objective optimization problem. In this case, each objective function corresponds to one task. In adversarial machine learning, attacks are traditionally measured on a single task using one objective function that measures the attack success rate (Croce and Hein 2020b) or the robustness (Tramer et al. 2020). However, in MTL when evaluating two or more attacks, there is no single metric, as there are multiple tasks and each task has its own objective function value. This makes the comparison of two different attacks in MTL challenging.
Therefore, we formulate the multi-task attack optimization problem with a multi-task–specific objective function that aligns with the standard practice of assessing model performance in MTL. A multi-task model’s performance is typically measured by Average Relative Accuracy (ARA), denoted as , as opposed to using absolute values (Sun et al. 2019; Zhang, Liu, and Guan 2022a). The ARA metric compares the performance of a given multi-task model to that of a baseline model :
@@
where represents the model loss difference for task before and after the attack to substitute the accuracy difference, since task loss in neural networks typically serves as a reliable indicator of task accuracy (i.e., higher task loss corresponds to lower task accuracy).
The optimization problem mentioned above can identify the optimal signed gradient direction vector but is intractable (Kreinovich, Lakeyev, and Noskov 1996; Horáček, Hladík, and Černỳ 2017). To address the problem, we apply the Taylor Expansion on , and reformulate it to an ILP problem as follows:

### v4

# Attacking All Tasks at Once Using Adversarial Examples in Multi-Task Learning

@@ Abstract
Visual content understanding frequently relies on multi-task models to extract robust representations of a single visual input for multiple downstream tasks. However, in comparison to extensively studied single-task models, the adversarial robustness of multi-task models has received significantly less attention and many questions remain unclear: 1) How robust are multi-task models to single task adversarial attacks, 2) Can adversarial attacks be designed to simultaneously attack all tasks in a multi-task model, and 3) How does parameter sharing across tasks affect multi-task model robustness to adversarial attacks? This paper aims to answer these questions through careful analysis and rigorous experimentation. First, we analyze the inherent drawbacks of two commonly-used adaptations of single-task white-box attacks in attacking multi-task models. We then propose a novel attack framework, Dynamic Gradient Balancing Attack (DGBA). Our framework poses the problem of attacking all tasks in a multi-task model as an optimization problem that can be efficiently solved through integer linear programming. Extensive evaluation on two popular MTL benchmarks, NYUv2 and Tiny-Taxonomy, demonstrates the effectiveness of DGBA compared to baselines in attacking both clean and adversarially trained multi-task models. Our results also reveal a fundamental trade-off between improving task accuracy via parameter sharing across tasks and undermining model robustness due to increased attack transferability from parameter sharing.
keywords:
[…]
Existing attempts on attacking multi-task models presents inherent drawbacks.
Formulate the MTL adversarial attack as an optimization problem.
Develop DGBA to efficiently solve the multi-task attack problem.
1 Introduction
Visual content understanding often employs multi-task learning (MTL) to tackle diverse downstream tasks such as image segmentation and depth estimation. MTL is favored for its high resource efficiency and task accuracy [ruder2017overview, zhang2022automtl, zhang2022tree, yao2020survey]. It involves constructing a unified multi-task model that extracts feature representations from visual inputs by enabling parameter sharing across tasks. As multi-task models are an inseparable component in many applications with high-security requirements, such as autonomous driving and robotics [leang2020dynamic, kokkinos2017ubernet, arcari2023bayesian], it is important to understand and assess their adversarial robustness to ensure the effective adoption in real applications.
@@ 1 Introduction
Machine learning models are susceptible to adversarial attacks where small, imperceptible perturbations to input data can lead to incorrect prediction [goodfellow2014explaining]. Although adversarial attacks have been extensively studied on single-task models [Delving, BARZ, xu2022securing], related work on multi-task models is scarce. A pioneering study [mao2020multitask] claimed that the adversarial robustness of deep neural networks (DNN) increases as the number of tasks increases when attacking the sum of the tasks’ loss, but its conclusion is overturned by a following study [ghamizi2022adversarial]. Multi-Task Attack [guo2020multi] tries to develop attacks in the MTL setting; however, the adversarial samples generated are specific to each task and thus fail to attack all tasks simultaneously. Some other related works [gurulingan2021uninet, sobh2021adversarial] start from generating adversarial examples by attacking one task at a time, which does not consider the multiple tasks simultaneously. UniNet [gurulingan2021uninet] goes one step further by evaluating a straightforward multi-task attack that targets the sum of all task losses; yet this naïve strategy still mainly degrades subset of tasks, leaving others comparatively robust. Motivated by these limitations, we systematically investigate two critical security research questions (RQ) for MTL in this work:
@@
RQ1: How robust are multi-task models to conventional single task adversarial attacks?
@@
RQ2: Can adversarial attacks be designed to simultaneously attack all tasks in a multi-task model?
To answer RQ1, we first categorize existing attempts on attacking multi-task models into two strategies, both relying on adapting existing single-task white-box attack approaches to multi-task models. We refer to these existing approaches as “adapted multi-task attacks” throughout the paper. We then analyze their inherent drawbacks that limit their attack effectiveness on multi-task models and motivate our proposed approach. (Section 2)
To answer RQ2, we propose a novel multi-task attack method, DGBA (Dynamic Gradient Balancing Multi-task Attack) to generate adversarial samples effective in attacking all tasks at once in a multi-task model. The central principle of DGBA is to harness established single-task attack techniques, thus avoiding redundant efforts, and at the same time, overcome the inherent limitations of existing adaptation strategies, thereby generating effective adversarial samples. The key insight is to dynamically balance gradients from multiple tasks in a multi-task model when creating adversarial samples that work across all tasks. (Section 3)
We conduct thorough experiments using common MTL benchmarks and branched multi-task models with different levels of parameter sharing. Experiments demonstrate that DGBA achieves up to 80.41% higher attack effectiveness on clean multi-task models and up to 18.65% higher attack effectiveness on adversarially trained multi-task models compared to baselines. DGBA remains the most effective attack approach in most cases regardless of the levels of parameter sharing and the attack strengths. (Section 4)
Our results in Section 4 also demonstrate that a higher degree of parameter sharing is associated with increased adversarial vulnerability of multi-task models, from which we raise the third research question RQ3: How does parameter sharing across tasks affect multi-task model robustness to adversarial attacks?. To answer RQ3, we empirically study the reason behind the tradeoff of parameter sharing and model robustness and find that improving a task’s accuracy and efficiency by sharing its parameters with other related tasks can increase the task’s vulnerability to adversarial attacks designated for these related tasks. This underlines the importance of balancing accuracy and robustness in multi-task model design. (Section 5)
We summarize our contributions as follows.
We categorize existing attempts on attacking multi-task models and analyze their inherent drawbacks.
We formulate the MTL adversarial attack as an optimization problem and develop DGBA to efficiently solve it.
We empirically evaluate the effectiveness of DGBA on multi-task models with various levels of parameter sharing and demonstrate that DGBA performs best for 7 out of 8 models on NYUv2 [silberman2012indoor] and 6 out of 8 on Tiny-Taskonomy [zamir2018taskonomy].
@@
We incorporate adversarial examples into the training of multi-task models to defend against adversarial attacks. The robustness of the models improves markedly, measured by the decreased task performance drop after attacking from to . When attacking these adversarially trained models, DGBA still outperforms baselines by up to .
We empirically demonstrate that parameter sharing can undermine model robustness due to increased attack transferability.
2 Background and Motivation
RQ1: How robust are multi-task models to conventional single task adversarial attacks? To answer this question, this section first gives background on existing white-box single-task attacks and our adversarial threat model. It then summarizes existing attempts on adapting single-task attacks to multi-task models and analyzes their drawbacks that motivate our proposed method DGBA. A detailed discussion of related work is in B.
2.1 Single Task White-Box Attacks
@@
In general, adversarial attacks can be formulated as follows [madry2018towards]. Let represent a clean input and its corresponding label. An attacker adds an adversarial perturbation to the input , to maximize the value of a loss function :
@@
where denotes the parameters of the trained model under attack, and represents the maximum amount the adversary can perturb the input according to a given -norm. For notational simplicity, we omit in our future derivations. Note that, although the input and the perturbation are matrices in vision tasks (defined at the pixel level), we keep matrix-valued variables in italic following the prevailing convention in adversarial-learning literature [croce2020reliable, xu2022securing]. A complete notation table is provided in A.
@@
Threat Model: In this paper, we focus on the untargeted white-box adversarial threat model [carlini2019evaluating] as this represents one of the strongest and most widely used adversarial machine learning formulations [carlini2017towards, dong2018boosting, croce2020reliable]. In this setup, the attacker has knowledge of the model structure, trained model parameters and the corresponding loss function . In terms of bounds on the adversarial perturbation, we use one of the most widely used norms, , in line with previous works [guo2020multi, xu2022securing, rathbun2022game].
@@
Single Task Attacks: In the white-box setting, one of the most prevalent strategies for generating the adversarial perturbation is to maximize the loss function by following the gradient ascent direction. This was originally done with the Fast Gradient Sign Method (FGSM) attack proposed in [goodfellow2014explaining]. Since the advent of FGSM, numerous improvements to the attack have been proposed. Although enumerating all the improvements in FGSM is beyond the scope of this paper, several important attack updates are worth noting. Updated attacks include the Projected Gradient Descent (PGD) attack [madry2018towards], which adds a randomized start and makes FGSM iterative. The Momentum Iterative Momentum (MIM) [dong2018boosting] adds momentum to the gradient ascent optimization. More recently, in APGD [croce2020reliable], an adaptive step size has been shown to be one of the most effective white-box attacks for a single task, even against adversarial trained models [mahmood2021robustness].
2.2 Adapting Single-Task Attacks to MTL
We first categorize existing methods on attacking multi-task models into two strategies, Single attack and Total attack. Both strategies rely on adapting single-task white-box attacks to multi-task models and thus are referred to as adapted multi-task attacks in this paper. We then analyze and empirically demonstrate their inherent flaws, which motivate the proposed multi-task attack method DGBA in Section 3.
@@ 2.2 Adapting Single-Task Attacks to MTL
In the case of a single task, untargeted white-box attack, an adversarial example can be generated [madry2018towards] iteratively:
@@
where is the current step, represents the perturbation function associated with a white-box attack, represents the loss function for a single task, is the magnitude of the perturbation added in the current iteration of the attack and . Lastly, is the projection operation [croce2020reliable] to bound the adversarial sample within a specified range.
In MTL, each input is associated with a set of true labels for tasks . Each task has its own task-specific loss function .
@@
Single Attack - The first way in which multi-task models can be attacked is by focusing on only a single task’s gradient and ignoring the gradients of the rest tasks. This approach is explored in [sobh2021adversarial, gurulingan2021uninet, li2023sibling]. For example, APGD [croce2020reliable] can be utilized to attack one of the task in the task set :
@@
where is a hyperparameter in APGD that controls the influence of previous update steps on the current update step. is the objective function of the task being attacked. We use Single-x to represent performing Single attack on a specific task x.
@@
Total Attack - The second way in which single task attacks can be converted to attack multi-task models is through totaling all associated task loss functions , via summation. This approach is used in [mao2020multitask, gurulingan2021uninet, ghamizi2022adversarial]. For example, single-task PGD [madry2018towards] can be adapted to Total-PGD:
@@
Limitations of Existing Attacks - Both the Single and Total attacks come with significant drawbacks. Single attack is designed to compromise one target task. It may incidentally affect the other tasks through adversarial transferability, but this multi-task effect is typically limited [mahmood2021robustness]. For example, in Figure 1, we show that when attacking the segmentation or the depth estimation task solely (Single-Segm or Single-Depth), the attack effectiveness on the normal prediction task is limited.
Likewise, the effectiveness of the Total attack is based on an underlying assumption: for an adversarial example to work across all tasks, no one task’s gradient should dominate to avoid the limited attack transferability issue that Single faces. We denote it as the non-dominant magnitude assumption. However, the issue of gradient dominance is well recognized within the MTL literature and has garnered significant attention in the field of MTL optimizers [yu2020gradient, navon2022multi]. Empirically, we observe from Figure 1 that the Total attack exhibits a pattern similar to Single-Segm, indicating that the segmentation task dominates the gradient directions.
We also consider a modified version of Total where we take the sign of the gradients before the summation. In this way, the non-dominant magnitude assumption can be circumvented. We denote this attack as SignTotal. However, we empirically show that this attack is also not effective in Section 4.2, as completely ignoring task’s gradient magnitudes also leads to a suboptimal attack.
@@
3 The Proposed Method: DGBA
RQ2: Can adversarial attacks be designed to simultaneously attack all tasks in a multi-task model (theoretically)? Dynamic Gradient Balancing Multi-task Attack (DGBA) builds on the success of existing single-task adversarial attacks, while addressing the challenge in attacking multi-task models. DGBA accomplishes this by dynamically balancing the gradients across tasks, to derive an adversarial perturbation that is effective on all tasks. In this section, we first formulate a new attack optimization problem tailored for multi-task models. Since the problem is intractable, DGBA reformulates it to an Integer Linear Programming (ILP) problem, and then generates adversarial samples by solving the ILP problem.
3.1 Multi-Task Attack Optimization
We first reformulate the original single-task adversarial optimization introduced in Eq. 1 by decomposing . Here, represents the magnitude of the perturbation. represents the signed gradient direction vector, with values .
The above formulation is used to attack a single task. However, attacking multiple tasks simultaneously in a multi-task model is fundamentally a multi-objective optimization problem. Specifically, since there are multiple tasks and each task has its own objective functions, there is no single metric that can measure the multi-task attack success as the attack success rate [croce2020reliable] used in single-task attack. Therefore, we formulate the multi-task attack optimization problem with a multi-task–specific objective function that aligns with the standard practice of assessing model performance in MTL.
A multi-task model’s performance is typically measured by Average Relative Accuracy (ARA), denoted as , as opposed to using absolute values [sun2019adashare, zhang2022automtl]. The ARA metric compares the performance of a given multi-task model to that of a baseline model :
@@ 3.1 Multi-Task Attack Optimization
where represents the model loss difference for task before and after the attack to substitute the accuracy difference, since task loss is widely used in MTL as an accuracy proxy, to indicate model quality (i.e., higher task loss corresponds to lower task accuracy) [sun2019adashare, zhang2022automtl, zhang2022tree]. Besides, considering the loss functions for vision tasks used in our work are Cross-Entropy Loss, Mean Squared Error, or Mean Absolute Error, we simply use instead of in the denominator.
The optimization problem in Eq. 7 can identify the optimal signed gradient direction vector that is effective on all tasks. However, this problem is intractable due to the high-dimensional, non-convex nature of the loss landscape across multiple tasks, where finding a global-optimal solution is computationally prohibitive [kreinovich1996approximate, horavcek2017interval]. Thus, we next explain a relaxed ILP formulation to solve the problem.
3.2 Optimization Solution via Relaxed LP
We reformulate Eq. 7 to an ILP problem by applying the Taylor Expansion on (see C for the detailed derivation):